Solve by factoring.
10t^2 - t - 2 = 0

Answer by lefty4ever26(59)

The trick to factoring these kinds of trinomials is to take the 1st constant(10) multiplied by the last constant (-2). This is -20. What two numbers multiplied together equal -20 but add together to equal the middle constant of your equation(-1). Well that would be -5 and 4. because -5*4=-20 and -5+4=-1
So to factor it we replace the -t with (-5t+4t) to get
10t^2-5t+4t-2=0 Then we need to factor by grouping
(10t^2-5t)+(4t-2)=0
5t(2t-1)+2(2t-1)=0 Then we combine to make the two factors
(5t+2)(2t-1)=0 Set each factor equal to 0 and solve for t.
5t+2=0 2t-1=0
t=-2/5 t=1/2
Hope this helps!